OpenStudy (rainbow_rocks03):

PLZ HELP ME I will medal+fan=you for BEST response. A bicycle sells for $240. The sales tax is 6%. What is the total cost of the bicycle? A. $240.06 B. $240.60 C. $246.00 D. $254.40

OpenStudy (studygurl14):

Do 240 + 240(6%) 240(6%) being the sales tax, and 240 being the cost of the bicyle itself.

OpenStudy (rainbow_rocks03):

240+240=480*6=2880

OpenStudy (studygurl14):

No...

OpenStudy (rainbow_rocks03):

oh

OpenStudy (studygurl14):

Do 240(6%) first. Remember that 6% = 0.06

OpenStudy (rainbow_rocks03):

240*0.06=14.4

OpenStudy (anonymous):

$254.4 You just multiply 0.06 with 240 and then add the answer to 240 :)

OpenStudy (rainbow_rocks03):

thx for the help @StudyGurl14 and @maggapoo

OpenStudy (studygurl14):

anytime. :)

OpenStudy (rainbow_rocks03):

:)